# The Theory of Economic Complexity

### v1

@@ 2 The Single Capability Model
We start with the basic model of economic complexity introduced in [1] and we will move later to other production functions.
This model assumes that an economy $c$ is endowed with capability $b$ with probability $r_{c,b}$ and that activity $p$ requires a capability $b$ with probability $q_{p,b}$. [^9] This deviates from previous work [1, 54] which tended to assume either a single $r$ and $q$ for all economies and activities or a known distribution.
For pedagogical reasons, we start with the case in which there is a single capability or factor, and an arbitrary number of economies and activities (that is $r_{c,b}\xrightarrow{}r_{c}$ and $q_{p,b}\xrightarrow{}q_{p}$). This case will allow us to get a basic intuition that we will then generalize to more complex functional forms. The advantage of starting with the single capability model is that it can be solved analytically, so it can provide us with a baseline understanding that we can extend to the numerical solutions derived for more complex models.
Let the output $Y_{cp}$ of economy $c$ in activity $p$ be given by the matrix:
@@
where $A$ is a constant or scale factor and $1-q_{p}(1-r_{c})$ is the probability that economy $c$ has the capability that product $p$ requires. This probability is written as a complement. That is, one minus the probability that the activity requires the capability ($q_{p}$) and the economy does not have it ($1-r_{c}$). In matrix form, the output matrix follows the form:
@@
Going forward, we always sort rows in descending order of $r$ and columns in ascending order of $q$. That is, the first cell of the matrix ($Y_{11}$) is the output of the economy with the highest probability of having the capability in the activity with the lowest probability of requiring it. This sorting convention will greatly facilitate the visual inspection of these matrices.
A key difference between this implementation of the model and previous work[1, 54] is that here we use the model to simulate an output matrix, whereas previous work used it to simulate a specialization matrix (which has already been through important manipulations and normalizations). We find that doing these normalization steps explicitly is essential for obtaining the right connection between the economic complexity index and the model parameters. So, next, we perform the manipulations applied to output matrices in the empirical literature to derive a metric of economic complexity. These are:
(i) Estimating the matrix of revealed comparative advantage or RCA $R_{cp}$ according to Balassa’s (1965) definition. This matrix normalizes the output matrix by the sum of its rows and columns and it is equivalent to a matrix comparing the observed output ($Y_{cp}$) with the expected output in a probabilistic model (see eqn. (3)). RCA is also known as the location quotient (LQ) in economic geography and innovation studies.
(ii) Estimating the binary specialization matrix $M_{cp}$. This is a matrix that is 1 if $R_{cp}\geq 1$ and 0 otherwise. This binary matrix is motivated in the empirical literature as a way to remove the tails of the $R_{cp}$ matrix, since the ratio definition of $R_{cp}$ results in larger variance for economies with low levels of output (small $Y_{c}=\sum_{p}Y_{cp}$) and activities with small markets (small $Y_{p}=\sum_{c}Y_{cp}$).
(iii) Estimating the $M_{cc^{\prime}}$ matrix. This is a square matrix connecting economies with similar specialization patterns and is the one used to derive the economic complexity index. This matrix is defined using the reciprocal averaging method known as the method of reflections [1], but it can also be defined as the product of a four matrices related to $M_{cp}$ (we will introduce the exact formula at that point).
@@
Since this is an inequality, we need to be careful about the signs of $(q_{p}-\langle q\rangle)$ and $(r_{c}-\langle r\rangle)$. Changes in sign flip the inequality operator. So what this condition means is that $R_{cp}\geq 1$ when $r_{c}\geq\langle r\rangle$ for $q_{p}-\langle q\rangle\geq 0$ and when $r_{c}<\langle r\rangle$ for $q_{p}-\langle q\rangle<0$. We can also get this condition intuitively by by considering the case when $q_{p}=\langle q\rangle$ or $r_{c}=\langle r\rangle$. In these two cases $R_{cp}=1$, meaning that these lines divide the matrix into regions where the values of $R_{cp}$ are higher or smaller than one. In sum, from the condition above $M_{cp}$ is a matrix divided into four quadrants:
$M_{cp}=1\quad\text{if}\quad r_{c}\geq\langle r\rangle\quad\&\quad q_{p}\geq\langle q\rangle$ (10)
$M_{cp}=0\quad\text{if}\quad r_{c}<\langle r\rangle\quad\&\quad q_{p}\geq\langle q\rangle$
$M_{cp}=0\quad\text{if}\quad r_{c}\geq\langle r\rangle\quad\&\quad q_{p}<\langle q\rangle$
$M_{cp}=1\quad\text{if}\quad r_{c}<\langle r\rangle\quad\&\quad q_{p}<\langle q\rangle$
This matrix represents a world in which countries with a high probability of having the capability ($r_{c}$ higher than average), specialize in products with high probability of requiring the capability ($q_{p}$ higher than average), and countries with low probability of having the capability specialize in products with low probability of requiring it. This is related to the idea of log super-modularity in trade theory [55].
@@
Estimating the first eigenvector of $M_{cc^{\prime}}$ is trivial because $M_{cc^{\prime}}$ is a stochastic matrix (each row adds to one). That means its first eigenvector will always be $\mathbf{1}$. This is easy to prove by summing $M_{cc^{\prime}}$ over $c^{\prime}$.
@@
We begin with the simplest case, which is an even number of economies and activities. We let also $\langle r\rangle$ and $\langle q\rangle$ be the medians of their distributions. In that case, $M_{cc^{\prime}}$ reduces to a block diagonal matrix with two blocks with values of $1/M_{p}$ (all economies have the same diversity and all activities the same ubiquity). That is:
@@
In this case, this eigenvector is also associated with the eigenvalue of one (this matrix is degenerate, meaning that it has more than one eigenvector associated with the same eigenvalue). [^11] In this case, all linear combinations of these eigenvectors are eigenvectors themselves. For example the vector $[a,a,b,b]$ is also an eigenvector, since we can construct it as a linear combination of $[1,1,1,1]$ and $[1,1,-1,-1]$ This eigenvector is easy to verify through multiplication.
@@
showing that in this example the second eigenvector of the $M_{cc^{\prime}}$ matrix or $ECI$ separate economies that are above or below average in their probability of having the only capability in the model. [^12] At this point, it is worth noting a key property of eigenvectors, which is that they have a freedom of sign. That is, if $e_{c}$ is an eigenvector of a matrix $M$ so is $-e_{c}$. This is trivial from the fact that if $Me_{c}=\lambda e_{c}$ then $M(-e_{c})=\lambda(-e_{c})$. This means that the eigenvector derivation of $ECI$ separates among economies based on their capability endowments, but is agnostic about which of the two clusters is the high-capability cluster. In the empirical literature, this is solved by iterating the system in eqs. 12 to estimate $ECI$ starting from an initial condition that is correlated with the high-capability cluster (e.g. initializing the system with diversity $M_{c}$) and stopping at an even iteration. Other methods to estimate complexity empirical (e.g. [13] also rely on an initial guess).
Figure 1: Graphical description of the four matrices involved in the single capability model for 10 countries and 20 products. In $cp$ matrices row represents economies (countries) and columns represent activities (products). Rows are sorted from highest $r_{c}$ to lowest $r_{c}$ and columns are sorted from lowest $q_{p}$ to highest $q_{p}$. That is, cell $(1,1)$ is the output of the country with the highest probability of having the capability on the product with the lowest probability of requiring it, and cell $(10,20)$ is the output of the country with the lowest probability of having a capability in the product with the highest probability of requiring it.
@@
From top left to bottom right, we start with the $Y_{cp}$ matrix, which the output matrix, the RCA matrix $R_{cp}$ which estimates the specialization of each economy in each activity, the binary specialization matrix $M_{cp}$, and the stochastic square matrix $M_{cc^{\prime}}$ from which we derive $ECI$. The output matrix $Y_{cp}$ shows a nested pattern. This is the tendency for the rows that are less filled to be subsets of the rows that are more filled. Nestedness is a well-known feature of matrices summarizing the geography of fine-grained economic activities, such as exports by country and product, employment by city and industry, or patents by city and technology [122]. This example shows how the matrix transformations simplify the complexity of the $Y_{cp}$ matrix, reducing it to a couple of clusters with above and below average probability of having a capability. But the high level of symmetry of this case limits our ability to explore more interesting properties of the method, such as the ability to separate capability endowments from simple measures of diversity. For that, we need to consider other cases.
Next, we focus on the case where the number of economies is odd and the number of activities is even (and where the averages of $r$ and $q$ are still their medians). For example, $N_{c}=5$ and $N_{p}=6$. This example is interesting, because unlike in the previous case where the diversity of economies and the ubiquity of activities was constant, in this case only the ubiquity of activities remains fixed. This example is important because it will teach us about the ability of $ECI$ to recover $r_{c}$, even when the most diverse economy is the one that has a probability of having a capability equal to the average ($r_{c}=\langle r\rangle$) (it is actually specialized in all activities).
@@
Figure 2: Graphical description of the four matrices involved in the single capability model for 11 economies (e.g. countries) and 20 activities (e.g products). In $cp$ matrices row represents economies (countries) and columns represent activities (products). Rows are sorted from highest $r_{c}$ to lowest $r_{c}$ and columns are sorted from lowest $q_{p}$ to highest $q_{p}$. That is, cell $(1,1)$ is the output of the country with the highest probability of having the capability on the product with the lowest probability of requiring it, and cell $(11,20)$ is the output of the country with the lowest probability of having a capability in the product with the highest probability of requiring it.
@@
$M_{cc^{\prime}}$ $=\frac{1}{M_{c}}(1+\frac{1}{N_{p}})\quad\text{if}\quad r_{c}\>\&\>r_{c^{\prime}}>\langle r\rangle\quad\textrm{or}\quad r_{c}\>\&\>r_{c^{\prime}}<\langle r\rangle$ (28)
$M_{cc^{\prime}}$ $=\frac{1}{M_{c}}(\frac{1}{N_{p}})\quad\text{if}\quad r_{c}\>>\langle r\rangle\quad\&\quad r_{c^{\prime}}<\langle r\rangle\ \text{and vice versa}$
$M_{cc^{\prime}}$ $=\frac{1}{N_{c}}(\frac{1}{N_{p}}+\frac{N_{c}-1}{M_{p}})\quad\text{if}\quad c\>=\>c^{\prime}\quad\&\quad r_{c}\>=\langle r\rangle$
$M_{cc^{\prime}}$ $=\frac{1}{N_{c}}(1+\frac{1}{N_{p}})\quad\text{if}\quad c\>\neq\>c^{\prime}\quad\&\quad r_{c}\>=\langle r\rangle$
$M_{cc^{\prime}}$ $=\frac{1}{M_{c}}(1+\frac{1}{N_{p}})\quad\text{if}\quad r_{c^{\prime}}\>=\langle r\rangle$
@@
$M_{cc^{\prime}}=\begin{bmatrix}\frac{1}{M_{c}}(\frac{N_{p}+1}{N_{p}})&\dots&\frac{1}{M_{c}}(\frac{N_{p}+1}{N_{p}})&\dots&\frac{1}{M_{c}}(\frac{1}{N_{p}})\\ \frac{1}{M_{c}}(\frac{N_{p}+1}{N_{p}}))&\dots&\frac{1}{M_{c}}(\frac{N_{p}+1}{N_{p}})&\dots&\frac{1}{M_{c}}(\frac{1}{N_{p}})\\ \frac{1}{N_{c}}(1+\frac{1}{N_{p}})&\dots&\frac{1}{N_{c}}(\frac{1}{N_{p}}+\frac{N_{c}-1}{M_{p}})&\dots&\frac{1}{N_{c}}(1+\frac{1}{N_{p}})\\ \frac{1}{M_{c}}(\frac{1}{N_{p}})&\dots&\frac{1}{M_{c}}(\frac{N_{p}+1}{N_{p}})&\dots&\frac{1}{M_{c}}(\frac{N_{p}+1}{N_{p}})\\ \frac{1}{M_{c}}(\frac{1}{N_{p}})&\dots&\frac{1}{M_{c}}(\frac{N_{p}+1}{N_{p}})&\dots&\frac{1}{M_{c}}(\frac{N_{p}+1}{N_{p}})\end{bmatrix}$ (29)
@@
Figure 3: Graphical description of the four matrices involved in the single capability model for 15 economies (e.g. countries) and 17 activities (e.g products). In $cp$ matrices row represents economies (countries) and columns represent activities (products). Rows are sorted from highest $r_{c}$ to lowest $r_{c}$ and columns are sorted from lowest $q_{p}$ to highest $q_{p}$. That is, cell $(1,1)$ is the output of the economy or country with the highest probability of having the capability on the product or activity with the lowest probability of requiring it, and cell $(11,20)$ is the output of the economy with the lowest probability of having a capability in the activity or product with the highest probability of requiring it.
@@
$M_{cp}v_{c}=\frac{a}{M_{c}}(1+\frac{1}{M_{p}})-\frac{a}{M_{c}}(\frac{1}{M_{p}})=\frac{a}{M_{c}}$ (32)
@@
$M_{cp}v_{c}=\frac{a}{M_{c}}(\frac{1}{M_{p}})-\frac{a}{M_{c}}(1+\frac{1}{M_{p}})=-\frac{a}{M_{c}}$ (33)
@@
Figure 3 presents these matrices in graphical form. We would like to notice two things about this version of the single capability model. The first one is that in this case $M_{cc^{\prime}}$ no longer has blocks of 00s. The second one is that this is also an example in which the highest diversity economy (the $8^{th}$ row in $M_{cp}$ is correctly identified as not being the economy with the highest probability of having the capability.
@@
In the next section we will use these figures to explore more complex forms of these model, involving multiple capabilities. We will then mode to different production functions to explore the generalizability of this result.

### v5

@@ 2 The Single Capability Model
We start with the basic model of economic complexity introduced numerically in [10]. This model assumes that an economy $c$ is endowed with capability $b$ with probability $r_{c,b}$ and that activity $p$ requires a capability $b$ with probability $q_{p,b}$. [^11] This deviates from previous work [10, 56] which tended to assume a distribution of $r$s and $q$s for all economies and activities instead of endowing each country and activity with an individual parameter.
For pedagogical reasons we start with the a single capability or factor and an arbitrary number of economies and activities (that is $r_{c,b}\xrightarrow{}r_{c}$ and $q_{p,b}\xrightarrow{}q_{p}$). This case will allow us to get a basic intuition that we will then generalize to more complex versions of the model. The advantage of starting with the single capability model is that we can derive its economic complexity index ($ECI$) analytically.
Let the output $Y_{cp}$ of economy $c$ in activity $p$ be given by the matrix [^12] We note that for $q_{p}=1$ this is the rectifier or ReLU function, which is a key activation function in neural networks.:
@@
where $A$ is a constant or scale factor and $1-q_{p}(1-r_{c})$ is the probability that economy $c$ has the capability that product $p$ requires. This probability is written as a complement. That is, one minus the probability that the activity requires a capability ($q_{p}$) that the economy does not have ($1-r_{c}$). In matrix form, the output matrix implied by this model is given by:
@@
Going forward, we sort rows in descending order of $r$ and columns in ascending order of $q$. That is, the first cell of the matrix ($Y_{11}$) is the output of the economy with the highest probability of having the capability in the activity with the lowest probability of requiring it. This sorting convention will greatly facilitate the visual inspection of these matrices.
A key difference between this implementation of the model and previous work[10, 56] is that here we use the model to simulate an output matrix ($Y_{cp}$), whereas previous work used it to simulate a specialization matrix (what we will later call $M_{cp}$). Specialization matrices have already been through important manipulations and normalizations. Our results show that these steps are essential for connecting $ECI$ with the model parameters. So, the remainder of this section will focus on performing the manipulations used in the empirical literature to this theoretical matrix. These manipulations are:
(i) Estimating the matrix of revealed comparative advantage or RCA $R_{cp}$ according to Balassa’s (1965) definition[125]. This matrix normalizes the output matrix $Y_{cp}$ by the sum of its rows and columns and it is equivalent to a matrix comparing the observed output ($Y_{cp}$) with the expected output in a probabilistic model (see eqn. (3)). RCA is also known as the location quotient (LQ) in economic geography and innovation studies.
(ii) Estimating the binary specialization matrix $M_{cp}$. This is a matrix that is 1 if $R_{cp}\geq 1$ and 0 otherwise. This binary matrix is motivated in the empirical literature as a means to remove the tails of the $R_{cp}$ matrix, since the ratio definition of $R_{cp}$ results in larger variance for economies with low levels of output (small $Y_{c}=\sum_{p}Y_{cp}$) and activities with small markets (small $Y_{p}=\sum_{c}Y_{cp}$).
(iii) Estimating the complexity matrix $M_{cc^{\prime}}$. This is a square matrix connecting economies with similar specialization patterns and is the one used to derive the economic complexity index. This matrix is defined using the reciprocal averaging method known as the method of reflections [10], but it can also be defined as the product of a four matrices (we will introduce the exact formula at that point).
@@
Since this is an inequality, we need to be careful about the signs of $(q_{p}-\langle q\rangle)$ and $(r_{c}-\langle r\rangle)$. Changes in sign flip the inequality operator. So what this condition means is that $R_{cp}\geq 1$ when $r_{c}\geq\langle r\rangle$ and $q_{p}-\langle q\rangle\geq 0$ or when $r_{c}<\langle r\rangle$ for $q_{p}-\langle q\rangle<0$. We can also get this condition intuitively by by considering the case when $q_{p}=\langle q\rangle$ or $r_{c}=\langle r\rangle$. In these two cases $R_{cp}=1$, meaning that these lines divide the matrix into regions where the values of $R_{cp}$ are higher or smaller than one. In sum, from the condition above $M_{cp}$ is a matrix divided into four quadrants:
$M_{cp}$ $=1\quad\text{if}\quad r_{c}\geq\langle r\rangle\quad\&\quad q_{p}\geq\langle q\rangle\quad$ (10)
$M_{cp}$ $=1\quad\text{if}\quad r_{c}<\langle r\rangle\quad\&\quad q_{p}<\langle q\rangle\quad$
$M_{cp}$ $=0\quad\text{if}\quad r_{c}<\langle r\rangle\quad\&\quad q_{p}\geq\langle q\rangle\quad$
$M_{cp}$ $=0\quad\text{if}\quad r_{c}\geq\langle r\rangle\quad\&\quad q_{p}<\langle q\rangle\quad\$
This matrix represents a world where countries with a high probability of having the capability ($r_{c}$ higher than average), specialize in products with high probability of requiring the capability ($q_{p}$ higher than average), and countries with low probability of having the capability specialize in products with low probability of requiring it. This is related to the idea of log super-modularity in trade theory [57].
@@
Estimating the first eigenvector of $M_{cc^{\prime}}$ is trivial because $M_{cc^{\prime}}$ is a stochastic matrix (each row adds to one). That means its first eigenvector will always be the vector $\mathbf{1}$. This is easy to prove by summing $M_{cc^{\prime}}$ over $c^{\prime}$.
@@
We begin with the simplest case, that of an even number of economies and activities. We let also $\langle r\rangle$ and $\langle q\rangle$ be the medians of their distributions. In that case, $M_{cc^{\prime}}$ reduces to a block diagonal matrix with two blocks with values of $1/M_{p}$ (all economies have the same diversity and all activities the same ubiquity). That is:
@@
In this case, this eigenvector is also associated with the eigenvalue of one (this matrix is degenerate, meaning that it has more than one eigenvector associated with the same eigenvalue). [^14] In this case, all linear combinations of these eigenvectors are eigenvectors themselves. For example the vector $[a,a,b,b]$ is also an eigenvector, since we can construct it as a linear combination of $[1,1,1,1]$ and [1,1,1,1][1,1,-1,-1]. This eigenvector is easy to verify through multiplication.
@@
showing that in this example the second eigenvector of the $M_{cc^{\prime}}$ matrix or $ECI$ separate economies that are above or below average in their probability of having the only capability in the model. [^15] At this point it is worth noting that a standard property of eigenvectors is that they have a freedom of sign. That is, if $e_{c}$ is an eigenvector of a matrix $M$ so is $e_{c}$. This is trivial from the fact that if $Me_{c}=\lambda e_{c}$ then $M(-e_{c})=\lambda(-e_{c})$. This means that the eigenvector derivation of $ECI$ separates among economies based on their capability endowments, but is agnostic about which of the two clusters is the high-capability cluster. In the empirical literature, this is solved by iterating the system of eqns. 12 to estimate $ECI$ starting from an initial condition that is correlated with the high-capability cluster (e.g. initializing the system with diversity $M_{c}$) and stopping at an even iteration. Other methods to estimate complexity empirical (e.g. [21] also rely on an initialization guess).
Figure 1: Graphical description of the four matrices involved in the single capability model for 10 countries and 20 products. In cpcp matrices row represents economies (countries) and columns represent activities (products). Rows are sorted from highest $r_{c}$ to lowest $r_{c}$ and columns are sorted from lowest $q_{p}$ to highest $q_{p}$. That is, cell $(1,1)$ is the output of the country with the highest probability of having the capability on the product with the lowest probability of requiring it, and cell $(10,20)$ is the output of the country with the lowest probability of having a capability in the product with the highest probability of requiring it.
@@
From top left to bottom right, we start with the output matrix ($Y_{cp}$), the specialization or RCA matrix ($R_{cp}$), the binary specialization matrix $M_{cp}$, and the complexity matrix $M_{cc^{\prime}}$ from which we derive $ECI$. The output matrix $Y_{cp}$ shows a nested pattern, which is a tendency for the rows that are less filled to be subsets of the rows that are more filled. Nestedness is a well-known feature of matrices summarizing the geography of fine-grained economic activities, such as exports by country and product, employment by city and industry, or patents by city and technology [126]. It is also a common feature of bipartite networks in ecology (e.g. pollinator networks or geographic specialization networks[127, 128]). This example shows how these transformations simplify $Y_{cp}$, reducing it to a couple of clusters with above and below average probability of having a capability. Yet, the symmetry of this example limits our ability to explore key properties of the method, such as the ability to separate capability endowments from simple measures of diversity. For that, we need to consider other cases.
Next, we focus on the case where the number of economies is odd and the number of activities is even (and where the averages of $r$ and $q$ are still their medians). For example, $N_{c}=5$ and $N_{p}=6$. This example is interesting, because unlike in the previous case where the diversity of economies and the ubiquity of activities was constant, here only the ubiquity of activities remains fixed. This example is important because it will teach us about the ability of $ECI$ to recover $r_{c}$, even when the most diverse economy is the one that has a probability of having a capability equal to the average ($r_{c}=\langle r\rangle$) (it is actually specialized in all activities).
@@
Figure 2: Graphical description of the four matrices involved in the single capability model for $1$ economies (e.g. countries) and 20 activities (e.g products). In cpcp matrices row represents economies (countries) and columns represent activities (products). Rows are sorted from highest $r_{c}$ to lowest $r_{c}$ and columns are sorted from lowest $q_{p}$ to highest $q_{p}$. That is, cell $(1,1)$ is the output of the country with the highest probability of having the capability on the product with the lowest probability of requiring it, and cell $(11,20)$ is the output of the country with the lowest probability of having a capability in the product with the highest probability of requiring it.
@@
$M_{cc^{\prime}}$ $=\frac{1}{M_{c}}\left(1+\frac{1}{N_{p}}\right)\quad\text{if}\quad r_{c}\>\&\>r_{c^{\prime}}>\langle r\rangle\quad\textrm{or}\quad r_{c}\>\&\>r_{c^{\prime}}<\langle r\rangle$ (28)
$M_{cc^{\prime}}$ $=\frac{1}{M_{c}N_{p}}\quad\text{if}\quad r_{c}\>>\langle r\rangle\quad\&\quad r_{c^{\prime}}<\langle r\rangle\ \text{and vice versa}$
$M_{cc^{\prime}}$ $=\frac{1}{N_{c}}\left(\frac{1}{N_{p}}+\frac{N_{c}-1}{M_{p}}\right)\quad\text{if}\quad c\>=\>c^{\prime}\quad\&\quad r_{c}\>=\langle r\rangle$
$M_{cc^{\prime}}$ $=\frac{1}{N_{c}}\left(1+\frac{1}{N_{p}}\right)\quad\text{if}\quad c\>\neq\>c^{\prime}\quad\&\quad r_{c}\>=\langle r\rangle$
$M_{cc^{\prime}}$ $=\frac{1}{M_{c}}\left(1+\frac{1}{N_{p}}\right)\quad\text{if}\quad r_{c^{\prime}}\>=\langle r\rangle$
@@
$M_{cc^{\prime}}=\begin{bmatrix}\frac{1}{M_{c}}\left(\frac{N_{p}+1}{N_{p}}\right)&\dots&\frac{1}{M_{c}}\left(\frac{N_{p}+1}{N_{p}}\right)&\dots&\frac{1}{M_{c}N_{p}}\\ \frac{1}{M_{c}}\left(\frac{N_{p}+1}{N_{p}}\right)&\dots&\frac{1}{M_{c}}\left(\frac{N_{p}+1}{N_{p}}\right)&\dots&\frac{1}{M_{c}N_{p}}\\ \frac{1}{N_{c}}\left(1+\frac{1}{N_{p}}\right)&\dots&\frac{1}{N_{c}}\left(\frac{1}{N_{p}}+\frac{N_{c}-1}{M_{p}}\right)&\dots&\frac{1}{N_{c}}\left(1+\frac{1}{N_{p}}\right)\\ \frac{1}{M_{c}}\left(\frac{1}{N_{p}}\right)&\dots&\frac{1}{M_{c}}\left(\frac{N_{p}+1}{N_{p}}\right)&\dots&\frac{1}{M_{c}}\left(\frac{N_{p}+1}{N_{p}}\right)\\ \frac{1}{M_{c}}(\frac{1}{N_{p}})&\dots&\frac{1}{M_{c}}(\frac{N_{p}+1}{N_{p}})&\dots&\frac{1}{M_{c}}(\frac{N_{p}+1}{N_{p}})\end{bmatrix}$ (29)
@@
Figure 3: Graphical description of the four matrices involved in the single capability model for 15 economies (e.g. countries) and 17 activities (e.g products). In cpcp matrices row represents economies (countries) and columns represent activities (products). Rows are sorted from highest $r_{c}$ to lowest $r_{c}$ and columns are sorted from lowest $q_{p}$ to highest $q_{p}$. That is, cell $(1,1)$ is the output of the economy or country with the highest probability of having the capability on the product or activity with the lowest probability of requiring it, and cell $(11,20)$ is the output of the economy with the lowest probability of having a capability in the activity or product with the highest probability of requiring it.
@@
$M_{cp}v_{c}=\frac{a}{M_{c}}\left(1+\frac{1}{M_{p}}\right)-\frac{a}{M_{c}}\frac{1}{M_{p}}=\frac{a}{M_{c}}$ (32)
@@
$M_{cp}v_{c}=\frac{a}{M_{c}}\left(\frac{1}{M_{p}}\right)-\frac{a}{M_{c}}\left(1+\frac{1}{M_{p}}\right)=-\frac{a}{M_{c}}$ (33)
@@
Figure 3 presents these matrices in graphical form. We would like to notice two things about this version of the single capability model. The first one is that in this case $M_{cc^{\prime}}$ no longer has blocks of 0s. The second one is that this is also an example in which the highest diversity economy (the $3^{rd}$ row in $M_{cp}$ is correctly identified as not being the economy with the highest probability of having the capability.
@@
In the next section we will use similar figures to explore more complex forms of this model, involving multiple and heterogeneous capabilities. We will then move to different production functions to explore the generalizability of this result.